# It’s Not the Target, It’s the Background: Rethinking Infrared Small Target Detection via Deep Patch-Free Low-Rank Representations

### v1

@@ I Introduction
Generally, an infrared image $\mathbf{D}$ can be decomposed into the following components [10]:
$\mathbf{D}=\mathbf{T}+\mathbf{B}+\mathbf{N},$ (1)
@@ III-B Compression and Reconstruction Module
Overview of the Proposed Compression and Reconstruction Module. Given an input feature map, we first apply a $3\times 3$ convolution to suppress local noise. The processed features are then passed through successive stages of the densely connected U-Net. At the lowest resolution level, a self-attention mechanism [4] is employed to capture long-range dependencies between features.
@@ IV-C7 The Effect of Pre-act ResBlock
TABLE V: Ablation Study on Model Size Using the IRSTD-1K Dataset.
@@
IV-C8 Model Size Discussion
Finally, we explore how the model’s architectural scale affects performance. We discuss the optimal number of channels and stages in our model. As shown in Tab. V, performance is poor when the number of channels is too small, due to insufficient filters to capture diverse image features. Conversely, simply increasing the number of channels can also degrade performance, as useful information becomes diluted within redundant channels. Regarding the number of stages, reducing stages causes out-of-memory (OOM) errors during training because the self-attention mechanism applied at the lowest resolution encounters exponentially larger feature maps, leading to quadratic complexity. On the other hand, increasing the number of stages excessively also hurts performance due to over-parameterization and resulting overfitting.
@@ IV-D Limitations
The limitations of our method can be summarized in the following three aspects:
Figure 6: Typical failure cases. The correlation between false alarm sources and the rest of the background is significantly weakened under strong noise interference.
First, a typical failure case is illustrated in Fig. 6. Under the interference of strong Gaussian white noise with a variance of 30, the correlation between false alarm sources at the feature level and the surrounding background is significantly weakened. Consequently, the proposed compress–reconstruct–subtract paradigm fails to effectively discriminate between true targets and false alarms. In such scenarios, learning target-specific features to facilitate further discrimination may be advantageous. We plan to investigate how to integrate target characteristic learning into the current framework in future work.
Second, most existing infrared small target detection methods rely on deep supervision [12] to achieve superior performance. However, conventional deep supervision strategies are ineffective in our current LRRNet due to the absence of well-defined constraints on intermediate outputs at different stages. A promising solution, inspired by DNANet [11], is to construct a feature pyramid by aggregating outputs from each encoder stage within the compress–reconstruction module, followed by fusion to generate a refined background feature map. This direction is left for future exploration.
Finally, our method only considers background translation invariance and neglects rotation invariance. An interesting observation is that current infrared small target detection algorithms exhibit performance degradation when dealing with rotated images [6]. In future work, we will explore the incorporation of rotation-equivariant convolutions [37] to address this challenge.
@@ V Conclusion
In this paper, we proposed LRRNet, an end-to-end network that directly models low-rank background priors in the image domain without relying on patch decomposition. By leveraging a compress–reconstruct–subtract paradigm, LRRNet effectively preserves small target features while reducing model complexity. Extensive experiments demonstrate that our method achieves superior accuracy, efficiency, and robustness compared to state-of-the-art approaches. This work highlights the promising potential of integrating interpretable low-rank priors for practical infrared small target detection. In future work, we intend to explore tailored solutions addressing the three aforementioned limitations within this paradigm, thereby further enhancing the framework’s capabilities.

### v2

@@ I Introduction
Generally, an infrared image $\mathbf{I}$ can be decomposed into the following components [10]:
$\mathbf{I}=\mathbf{T}+\mathbf{B}+\mathbf{N},$ (1)
@@ III-B Compression and Reconstruction Module
Overview of the Proposed Compression and Reconstruction Module. The proposed compression-reconstruction module can be viewed as a specialized mapping $\mathcal{F}:\mathbf{F}_{I}\rightarrow\mathbf{F}_{B}$, where $\mathbf{F}_{I}\in\mathbb{R}^{W\times H\times C}$ denotes the feature map extracted from the original image, and $\mathbf{F}_{B}$ represents the corresponding reconstructed background feature map. Notably, $\mathbf{F}_{B}$ shares the same spatial and channel dimensions as $\mathbf{F}_{I}$. Given an input feature map, we first apply a $3\times 3$ convolution to suppress local noise. The processed features are then passed through successive stages of the densely connected U-Net. At the lowest resolution level, a self-attention mechanism [4] is employed to capture long-range dependencies between features.
@@
At the lowest resolution level, a standard query-key-value (qkv) self-attention mechanism [47] is applied. Since the feature map $\mathbf{X}$ has undergone four downsampling operations, its spatial resolution is reduced from the original $H\times W$ to $\frac{H}{16}\times\frac{W}{16}$ resulting in negligible computational overhead. The standard qkv self-attention is formulated as:
$\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{Softmax}\left(\frac{\mathbf{QK}^{\top}}{\sqrt{d_{k}}}\right)\mathbf{V},$ (4)
where $\mathbf{Q}=\mathbf{X}\mathbf{W}^{Q}$, $\mathbf{K}=\mathbf{X}\mathbf{W}^{K}$, and $\mathbf{V}=\mathbf{X}\mathbf{W}^{V}$ are the query, key, and value projections of the input feature map $\mathbf{X}$, and $d_{k}$ denotes the dimensionality of the key vectors used for scaling. Recent studies [48] have shown that when the number of channels is large, the information captured across different channels tends to be highly redundant. As a result, employing multi-head attention in such scenarios provides limited benefit. In our design, self-attention is applied at the lowest resolution level, where the channel dimension is the largest within the network. Therefore, we adopt single-head attention instead of a multi-head formulation to reduce redundancy and computational overhead.
@@ IV-C7 The Effect of Pre-act ResBlock
TABLE V: Ablation Study on Hyperparameter Using the IRSTD-1K Dataset.
Hyperparameter Discussion IRSTD-1K
[…]
The effect of the number of $\lambda$ in loss function
$\lambda$ = 0.0 70.84 67.23 91.92 12.53
$\lambda$ = 0.1 72.35 69.51 95.96 3.36
$\lambda$ = 0.2 72.05 68.43 92.26 10.36
$\lambda$ = 0.3 70.86 69.17 90.57 8.67
$\lambda$ = 0.4 71.41 69.20 92.26 13.97
IV-C8 Hyperparameter Discussion
The key hyperparameter in our LRRNet is the weight $\lambda$ in the loss function, which balances the optimization contributions from the segmentation and reconstruction branches. The impact of different $\lambda$ values on the final performance is shown in Tab. V. When $\lambda=0$, the model lacks constraints from the compression-reconstruction process, leading to degraded performance. On the other hand, a large $\lambda$ causes the reconstruction branch to dominate the optimization. An interesting observation during training is that when $\lambda$ is large, the segmentation loss remains nearly stagnant in the early stages and only begins to decrease when the reconstruction loss has been sufficiently minimized.
TABLE VI: Ablation Study on Model Size Using the IRSTD-1K Dataset.
@@ IV-C8 Hyperparameter Discussion
The effect of the number of ResBlocks in DB
number = 1 69.73 68.30 94.91 35.51
number = 2 72.35 69.51 95.96 3.36
number = 3 69.23 68.55 93.27 49.71
IV-C9 Model Size Discussion
Finally, we investigate the impact of architectural scale on model performance, focusing on the optimal configuration of channels, stages, and the number of ResBlocks per stage. As shown in Tab. VI, performance degrades when the number of channels is too small, as insufficient filters hinder the model’s ability to capture diverse image features. Conversely, blindly increasing the number of channels may also impair performance, as the additional capacity can lead to feature redundancy and diluted useful information. In terms of stage design, reducing the number of stages leads to out-of-memory (OOM) errors during training. This is because the self-attention module, applied at the lowest-resolution feature map, incurs a quadratic computational cost, which becomes prohibitive as the resolution increases. On the other hand, excessively increasing the number of stages results in over-parameterization, making the model prone to overfitting and degrading generalization. As illustrated in Fig. 3, due to the densely connected design of LRRNet, each decoder stage contains one more ResBlock than its corresponding encoder stage. For clarity, Tab. VI only reports the number of ResBlocks in each encoder stage under different configurations. It can be observed that the model achieves the best performance when each encoder stage contains two ResBlocks.
@@ IV-D Limitations
The limitations of our method can be summarized in the following two aspects:
Figure 6: Typical failure cases. The correlation between false alarm sources and the surrounding background is significantly diminished under strong noise interference.
First, a representative failure case is depicted in Fig. 6. When subjected to strong Gaussian white noise with a variance of 30, the correlation at the feature level between false alarm sources and the surrounding background becomes significantly attenuated. As a result, the proposed compress–reconstruct–subtract paradigm is unable to effectively distinguish true targets from false alarms. In such challenging scenarios, incorporating target-specific feature learning may provide enhanced discrimination capability. We intend to explore the integration of target-aware representations into our current framework in future work.
Second, most state-of-the-art infrared small target detection approaches leverage deep supervision [12] to achieve high detection accuracy. However, traditional deep supervision schemes are not directly applicable to our LRRNet due to the lack of explicit constraints on intermediate outputs across different stages. Inspired by DNANet [11], a potential improvement is to construct a hierarchical feature pyramid by aggregating multi-scale outputs from each encoder stage within the compress–reconstruct module, followed by feature fusion to produce a more refined background representation. We leave this direction for future investigation.
@@ V Conclusion
In this paper, we proposed LRRNet, an end-to-end network that directly models low-rank background priors in the image domain without relying on patch decomposition. By leveraging a compress–reconstruct–subtract paradigm, LRRNet effectively preserves small target features while reducing model complexity. Extensive experiments demonstrate that our method achieves superior accuracy, efficiency, and robustness compared to state-of-the-art approaches. This work highlights the promising potential of integrating interpretable low-rank priors for practical infrared small target detection. In future work, we intend to explore tailored solutions addressing the two aforementioned limitations within this paradigm, thereby further enhancing the framework’s capabilities.